OpenStudy (anonymous):

Integral of (3x+1)/(x^3(3x+2)(x-1))

myininaya (myininaya):

partial fractions is needed here

myininaya (myininaya):

hint \[\frac{A}{x}+\frac{B}{x^2}+\frac{C}{x^3}+\frac{D}{3x+2}+\frac{E}{x-1}\]

OpenStudy (anonymous):

What do you do with all of that? I see that, in my notes, it adds up all of that using cross multiplication. But how do you do that with 5 fractions?

myininaya (myininaya):

it is suppose to equal \[\frac{3x+1}{x^3(3x+2)(x-1)}\] to find when that sum of fractions equal the one in your integral you need to find a common denominator so you can write the fractions as one

myininaya (myininaya):

\[\frac{A}{x} \cdot \frac{x^2(3x+2)(x-1)}{x^2(3x+2)(x-1)}\] for example the first fraction needs to be multiplied by this 1 can you do the others?

OpenStudy (anonymous):

Thank you, I think I got it from here!

myininaya (myininaya):

Ok it isn't that bad I know it seems a little long and maybe it would be too long if it was a test (maybe)

OpenStudy (anonymous):

Its homework... this teacher has a knack for making things look super complicated. Thanks a bunch